A 1μC charge is keeping a 1kg mass that is 1 m away from it from falling. Find the amount of charge that this mass is carrying. Assume point charges.
Answer #1

let
q1 = 1 micro C
m = 1 kg
d = 1 m
let q2 is the charge of 1 kg mass.

in the equilibrium, Fe = Fg

k*q1*q2/d^2 = m*g

q2 = m*g*d^2/(k*q1)

= 1*9.8*1^2/(9*10^9*1*10^-6)

= 0.00109 C (or) 1090 micro C <<<<<<<<-----Answer